Hi how is everyone today? Hope someone can help me this problem.
Science and medicine. Martina leaves home at 9 A.M., bicycling at a rate of
20 mi/h. Two hours later, John leaves, driving at the rate of 40 mi/h. At what time will John catch up with Martina? Show your work.

Answer by checkley77(12569)

20T=40(T-2)
20T=40T-80
20T-40T=-80
-20T=-80
T=-80/-20
T=4 HOURS AFTER MARTINA LEAVES THEY WILL MEET. OR @ 1:00 P.M.
PROOF:
20*4=40(4-2)
80=40*2
80=80